frank went 16 miles at one speed and then caame back going 4mph faster. if the return trip took 40mins less time find the 2 speeds.
Answer by mananth(16946)

Forward time - return time = 40 minutes = 2/3 hours
Let forward rate be x
return rate = x+4
16/x - 16/(x+4) = 2/3
LCD
3x(x+4)
48(x+4) -48x = 2*3x(x+4)
48x + 192 -48x = 6x^2+24x
6x^22 +24x -192=0
/6
x^2+4x-32=0
x^2+8x-4x-32=0
x(x+8)-4(x+8) =0
(x+8)(x-4) =0
x=4 the positive value
4 mph & 8 mph